# Random Access for LEO Satellite Communication Systems via Deep Learning

### v1

@@ V-B Random Access Performance
This subsection presents the end-to-end random access performance of the proposed framework. To ensure a fair comparison across different schemes, all approaches were configured with the same number of available preambles, thereby maintaining an equal number of scheduled PUSCH resources. However, in the case of DCRA, additional PUSCH resources are allocated to certain users when a preamble collision is detected. More specifically, if two users select the same preamble, DCRA assigns two separate PUSCH resources to that preamble to resolve the collision. As a result, even with the same number of available preambles, the total number of scheduled PUSCH resources in DCRA differs, since multiple grants may be allocated for a single preamble. For the S-eCD scheme, both the ideal case with zero RTT uncertainty and the practical case with the same RTT uncertainty setting as other methods were evaluated. Unless otherwise specified, all results are obtained under the configuration of 8 receive antennas, the TDL-D channel model and an SNR of 10 dB. Following the specification in [TWC23_DLDouble], the DCRA scheme uses $K=3$, whereas the proposed framework adopts $K=6$.
@@
Fig. 6 illustrates the average random access delay for each scheme. Here, delay is defined as the time elapsed until a user either successfully completes the random access procedure or exhausts the maximum number of retrial attempts. Among the compared schemes, the practical S-eCD exhibits the longest delay due to its poor preamble collision detection accuracy, which prevents users from identifying collisions at an early stage and thus delays contention resolution. In contrast, the ideal S-eCD enables earlier backoff decisions through moderate collision detection performance, resulting in a relatively shorter delay. The performance of the proposed framework and DCRA varies depending on the user density. In low-density environments, the proposed framework achieves shorter access delays. This is because it allows collided users to probabilistically back off after Step 2 via the opportunistic transmission scheme, whereas DCRA mandates unconditional Step 3 transmissions using additional PUSCH resources even when collisions are detected. However, in high-density scenarios, DCRA becomes more efficient. As the probability of three or more users selecting the same preamble increases, most collided users in DCRA are forced to back off at Step 2, effectively avoiding unnecessary Step 3 transmissions. On the other hand, the proposed framework still permits Step 3 transmissions, which can increase delay under heavy contention. This behavior results in a crossing point where the delay of DCRA becomes lower than that of the proposed framework as the number of active users increases.
@@
Fig. 7 and Fig. 8 show the average number of successful random access users and the PUSCH utilization, respectively. In general, if the number of successful users increases linearly with the number of active users, it indicates that most active users successfully complete the random access process which is desirable. The proposed framework demonstrates superior performance in terms of the average number of successful users by not only detecting preamble collisions effectively but also increasing the chance of successful access through its opportunistic transmission scheme. This advantage is also reflected in Fig. 8, where the proposed framework consistently achieves higher PUSCH utilization compared to the other schemes. In contrast, the PUSCH utilization of DCRA is relatively low because it allocates two PUSCH resources to the colliding users, thereby inducing additional contention. S-eCD shows moderate performance in the ideal case but performs poorly in the practical case, primarily because it does not account for channel delay spread and RTT uncertainty—factors that are present in realistic satellite environments. Moreover, even in the ideal case, S-eCD quickly saturates and then exhibits rapid performance degradation, indicating low RACH resource efficiency. In summary, the proposed framework demonstrates the highest overall resource efficiency under identical RACH and PUSCH configurations, outperforming other baseline schemes across a wide range of user loads. To clearly demonstrate the superiority of the proposed framework, Table IV summarizes the performance of the proposed method and the comparison schemes across key performance metrics.
@@
TABLE IV: Summary of random access performance metrics under different user densities.
Performance Metric Scheme Low Density High Density
Average number of successful users Proposed Very High High
DCRA [TWC23_DLDouble] High High
Ideal S-eCD [TVT21_SeCD] Very High Low
Practical S-eCD [TVT21_SeCD] Low Very Low
PUSCH utilization Proposed Very High High
DCRA [TWC23_DLDouble] High High
Ideal S-eCD [TVT21_SeCD] Very High Low
Practical S-eCD [TVT21_SeCD] Very Low Very Low
Average random access delay Proposed Very Low Low
DCRA [TWC23_DLDouble] Very Low Very Low
Ideal S-eCD [TVT21_SeCD] Low High
Practical S-eCD [TVT21_SeCD] Very High Very High
@@ VI Conclusion
Future research may extend this work by exploring adaptive random access strategies that account for user mobility and traffic heterogeneity. In particular, developing robust and scalable Step 3 transmission control mechanisms based on spatiotemporal prediction would be a promising direction.

### v2

@@ V-B Random Access Performance
This subsection presents the end-to-end random access performance of the proposed framework. To ensure a fair comparison across different schemes, all approaches were configured with the same number of available preambles, thereby maintaining an equal number of scheduled PUSCH resources. However, in the case of DCRA, additional PUSCH resources are allocated to certain users when a preamble collision is detected. More specifically, if two users select the same preamble, DCRA assigns two separate PUSCH resources to that preamble to resolve the collision. As a result, even with the same number of available preambles, the total number of scheduled PUSCH resources in DCRA differs, since multiple grants may be allocated for a single preamble. For the S-eCD scheme, both the ideal case with zero RTT uncertainty and the practical case with the same RTT uncertainty setting as other methods were evaluated. Unless otherwise specified, all results are obtained under the configuration of 8 receive antennas, the TDL-D channel model and an SNR of 10 dB. Following the specification in [43], the DCRA scheme uses $K=3$, whereas the proposed framework adopts $K=6$. The number of active users is set to a maximum of 52 by accounting for the LEO SatCom beam footprint size and the users’ access-attempt interval, which together determine the plausible number of simultaneous access attempts within a beam [5, 6, 26].
Fig. 7 illustrates the average random access delay for each scheme, defined as the time elapsed until a user successfully completes the random access procedure or exhausts the maximum number of retrials. This delay is averaged over mixed realizations that naturally include both collision-free and collision-present cases under the considered traffic and channel settings. The practical S-eCD yields the longest delay due to limited early collision detection capability, whereas the ideal S-eCD achieves shorter delay through earlier backoff decisions.
@@
DRA and MT-eCD exhibit short delays at low user densities because they enforce immediate backoff upon collision detection, but the delay increases rapidly under heavy contention due to repeated deferrals. Compared with DCRA, the proposed framework achieves lower delay in low-density regimes by allowing collided users to probabilistically defer Step 3 after Step 2, while DCRA mandates Step 3 transmissions with additional PUSCH grants. As the user density increases, higher-order collisions become more frequent, and the two schemes exhibit different trade-offs between preserving access opportunities for collided users and suppressing repeated transmissions, which can lead to a crossing point in delay performance. For reference, we also include the conventional 3GPP random access baseline, confirming that early preamble collision detection does not incur a delay penalty in the considered operating range.
@@
Fig. 8 shows the average number of successful random access users. In general, when the number of successful users increases nearly linearly with the number of active users, it implies that most users successfully complete the random access procedure, which is desirable. From this perspective, schemes that rely on immediate backoff upon collision detection, without additional contention-resolution actions, tend to suffer a sharp performance degradation as the user density increases. In contrast, both the proposed framework and DCRA sustain higher success performance by leveraging early collision awareness and resolving contention beyond simple backoff decisions. In particular, the proposed framework tends to achieve a higher number of successful accesses in the low-to-high density regime, because its collision-aware opportunistic Step 3 transmission provides additional access opportunities for collided users without deterministically discarding all of them.
The superiority of the proposed framework over DCRA is also evident in terms of PUSCH utilization. PUSCH utilization is defined as the ratio of the Step 3 PUSCH resources that are successfully used without collision to the total PUSCH resources allocated for Step 3. As shown in Fig. 9, the proposed framework achieves higher PUSCH utilization than DCRA. This is because, in DCRA, when two users select the same preamble in Step 1, the base station allocates additional Step 3 PUSCH resources, resulting in two PUSCH resources being associated with a single preamble and enabling contention-based transmissions in Step 3. If the two users transmit on different PUSCH resources, no PUSCH resource is wasted; however, if both users attempt Step 3 on the same PUSCH resource, then the two allocated PUSCH resources are effectively wasted due to the collision. In contrast, the proposed framework does not allocate additional Step 3 PUSCH resources. Instead, users attempt Step 3 opportunistically based on the collision information, which reduces unnecessary PUSCH allocations and yields higher resource efficiency.
@@
Fig. 10 investigates the impact of $K$ (the maximum collision multiplicity considered by the classifier) on the proposed framework. As $K$ increases, the framework can differentiate a wider range of collision multiplicities and apply more fine-grained Step 3 attempt control, which increases the number of successful users and delays the onset of performance saturation as the user density grows. Consistently, the random access delay is reduced over a wider user-density range because more collision cases can be handled without immediately forcing retransmission/backoff. However, the performance gain from increasing $K$ eventually saturates, indicating diminishing returns beyond a certain $K$ while the classification task becomes more complex due to the increased number of collision scenarios.
Figure 10: Performance of the proposed framework versus the number of active users under different $K$: (a) average number of successful users and (b) average random access delay.
Fig. 11 evaluates the proposed framework in a lower-SNR regime, $\text{SNR}\in[-17,-15,-13]$ dB, to assess robustness. As the SNR decreases, the average number of successful users decreases and the average random access delay increases, since collision signatures in the correlation sequence become less distinguishable and misclassification becomes more likely. Nevertheless, the proposed framework shows consistent trends with respect to user density across SNRs, indicating robust behavior rather than unstable performance. These results also serve as a proxy for large-scale attenuation effects (e.g., rain attenuation and shadowing), which mainly shift the effective operating SNR without fundamentally changing the collision-induced correlation structures.
Figure 11: Performance of the proposed framework versus the number of active users under different SNRs: (a) average number of successful users and (b) average random access delay.
Fig. 12 examines the sensitivity of the proposed framework to residual Doppler (residual frequency offset) after Doppler compensation/tracking. According to 3GPP assumptions for frequency error after compensation, the residual offset can be on the order of 10 Hz [4] however, to stress-test the proposed classifier and avoid optimistic conclusions, we additionally consider a much harsher case with a maximum residual Doppler of 250 Hz. Increased residual Doppler introduces phase rotation across the correlation processing interval, which blurs the collision-dependent structures and reduces the reliability of collision-multiplicity estimation. As a result, fewer users complete random access successfully and the access delay increases due to additional contention-resolution and retransmission events. Despite this stress-test setting, the degradation remains limited and the user-density-dependent trends are preserved, supporting the robustness of the proposed framework against residual Doppler.
Figure 12: Performance of the proposed framework versus the number of active users under residual Doppler: (a) average number of successful users and (b) average random access delay.
Overall, these results demonstrate that the proposed framework effectively improves end-to-end random access performance by leveraging early collision awareness and collision-multiplicity-guided Step 3 control. Across a wide range of user densities, it provides a favorable success–delay trade-off compared with the considered baselines, while maintaining robust performance trends under different design and impairment conditions, including the choice of $K$, low-SNR regimes, and residual Doppler.
@@ VI Conclusion
Future research may extend this work by incorporating inter-satellite interference under more aggressive random access resource reuse across satellites/beams. Another promising direction is to integrate PHY-layer collision-type estimation with learning-based (e.g., RL) access-control policies for cross-layer contention management under dynamic traffic and coverage conditions.